Exercise 5-16 a,c (Video)

Oriole Company estimates that variable costs will be 60.00% of sales, and fixed costs will total $600,000. The selling price of the product is $4.

Compute the break-even point in (1) units and (2) dollars.
(1) Break-even sales units
(2) Break-even sales $

LINK TO TEXT

LINK TO TEXT

Assuming actual sales are $2,000,000, compute the margin of safety in (1) dollars and (2) as a ratio.
(1) Margin of safety $
(2) Margin of safety ratio %
Click if you would like to Show Work for this question:

Open Show Work
Answer #1

Contribution Margin per unit = 100% - Variable cost ratio = 100% - 60% = 40% of sales

= 40% * 4 = 1.60

Breakeven sales units = Fixed cost/Contribution margin per unit

= 600,000/1.60 = 375,000 units

Breakeven sales = Fixed cost/Contribution margin ratio

= 600,000/40% = 1,500,000

Margin of Safety = Sales - Breakeven sales

= 2,000,000 - 1,500,000 = 500,000

Margin of Safety ratio = Margin of Safety/Sales

= 500,000/2,000,000 = 25%